OpenStudy (anonymous):

what Yp should i use if i have this eqn? y''-y'-6y=1+e^-2x

ganeshie8 (ganeshie8):

try \(Y_p = A + Be^{-2x}\)

OpenStudy (anonymous):

why not Ax+B ?

OpenStudy (anonymous):

isit because 1 on the RHS doesnt have an x counter part?

ganeshie8 (ganeshie8):

yes but with that guess, the exponent is disappearing on left hand side

ganeshie8 (ganeshie8):

@experimentX

OpenStudy (anonymous):

i used the eqn you give but i cant find the coefficients.. i used Ax+B and i got yp= -1/6+e^-2x

OpenStudy (experimentx):

i gotta go out soon ... what method are you using?

OpenStudy (anonymous):

undertermined coefficients

ganeshie8 (ganeshie8):

\(y= A + Be^{-2x}\) \(y'= - 2Be^{-2x}\) \(y''= 4Be^{-2x}\) -------------------- \(y'' - y' - 6y = -6A = 1 + e^{-2x}\) \(\implies A = \dfrac{-1}{6}\)

ganeshie8 (ganeshie8):

B just disappeared... :/

ganeshie8 (ganeshie8):

should i change the initial guess @experimentX

OpenStudy (experimentx):

try individually ... first for Y_p = x Y_p = e^(2x) or what ever .... then use superposition theorem.

OpenStudy (experimentx):

that you are using A is giving wrong answer.

OpenStudy (experimentx):

it must be Cx+D

OpenStudy (experimentx):

or try y = Ax+B + Ce^( .... x) good luck!! gotta go out for a while. probably won't be back.

OpenStudy (experimentx):

http://en.wikipedia.org/wiki/Method_of_undetermined_coefficients#Typical_forms_of_the_particular_integral the superposition of first and second.

ganeshie8 (ganeshie8):

okay... wil try thnks @experimentX :)

ganeshie8 (ganeshie8):

\(\dfrac{-1}{6} - \dfrac{1}{5}e^{-2x}\)

ganeshie8 (ganeshie8):

wolfram is giving that^^

OpenStudy (anonymous):

so that means Ax+B + ce^-2x is correct?

OpenStudy (experimentx):

okay okay ... this is 1/6 http://www.wolframalpha.com/input/?i=y%27%27-y%27-6y%3D1

OpenStudy (experimentx):

the other is http://www.wolframalpha.com/input/?i=y%27%27-y%27-6y%3De%5E%28-2x%29

OpenStudy (anonymous):

hmm.. does this mean it cant be solved by undetermined coefficients? do method power series apply on this?

ganeshie8 (ganeshie8):

\(Y_p = A + Bxe^{-2x}\)

ganeshie8 (ganeshie8):

since the first guess \(Y_p = A + Be^{-2x}\) didnt work, we make a new guess \(Y_p = A + Bxe^{-2x}\)

OpenStudy (anonymous):

okay ill try

OpenStudy (anonymous):

yeah! i got it! yay! xD

OpenStudy (anonymous):

thanks! but, i'm not really sure how to guess what's the right Yp.. is there some kind of a clue?

ganeshie8 (ganeshie8):

\(y = A + Bxe^{-2x}\) \(y' = Be^{-2x} - 2Bxe^{-2x}\) \(y'' = -2Be^{-2x} - 2Be^{-2x} + 4Bxe^{-2x}\) ----------------------------------------------- \(y'' - y' - 6y = 6Bxe^{-2x} - 5Be^{-2x} - 6A - 6Bxe^{-2x} = 1 + e^{-2x}\)

ganeshie8 (ganeshie8):

\(y'' - y' - 6y = - 5Be^{-2x} - 6A = 1 + e^{-2x}\) \(A = \dfrac{-1}{6}\) \(B = \dfrac{-1}{5}\)

OpenStudy (anonymous):

yup! i got -1/6 and -1/5 too! :D

ganeshie8 (ganeshie8):

it is the right solution as plugging it in satisfies the right hand side, right ?

ganeshie8 (ganeshie8):

oh you're asking how to make initial guess for \(Y_p\) ?

OpenStudy (anonymous):

yeah.. i'm not really sure whenever i guess..

OpenStudy (anonymous):

my prof have this solution but im not really sure how he does it

OpenStudy (anonymous):

how does he know that it precicely the solution for the complementary equation?

ganeshie8 (ganeshie8):

you fine wid solving the complementary solution, right ?

OpenStudy (anonymous):

complementary means the chaaracteristic equation?

ganeshie8 (ganeshie8):

yup solution from characteristic equaiton

ganeshie8 (ganeshie8):

\(y'' + 2y' + y = xe^{-x}\) complementary solution : \(y_c(x) = c_1e^{-x} + c_2xe^{-x}\)

ganeshie8 (ganeshie8):

complementary solution the solution for \(y'' + 2y' + y = 0\)

OpenStudy (anonymous):

yeah because it has 2 distinct roots

OpenStudy (anonymous):

ohh equal

ganeshie8 (ganeshie8):

\(y'' + 2y' + y = xe^{-x}\) \(y_c(x) = c_1e^{-x} + c_2xe^{-x}\) To find the particular solution, look at right hand side and make a guess

ganeshie8 (ganeshie8):

\(y'' + 2y' + y =\color{red}{ xe^{-x}}\)

ganeshie8 (ganeshie8):

the guess for polynomial \(\color{red}{x}\): \(Ax + B\) the guess for exponent \(\color{red}{e^{-x}}\) : \(Ce^{-x}\) so the overall guess is : \((Ax+B)(Ce^{-x})\)

ganeshie8 (ganeshie8):

but C is just an unnecessary constant, so so the overall guess is : \((Ax+B)e^{-x}\)

ganeshie8 (ganeshie8):

first guess : \(Y_p = (Ax+B)e^{-x}\)

ganeshie8 (ganeshie8):

fine till here ?

OpenStudy (anonymous):

yup! :D thats my guess too.. but why it isnt right?

ganeshie8 (ganeshie8):

it is right, and it is indeed a solution. but the bad thing is it is `part of ur complementary solution family`. it is not constrained to the `right hand side specific function`.

ganeshie8 (ganeshie8):

\(y_c(x) = c_1e^{-x} + c_2xe^{-x}\) plugin \(c_1 = B, ~ c_2 = A\) above^

ganeshie8 (ganeshie8):

you get back ur first guess, right ?

ganeshie8 (ganeshie8):

so ur `first guess` is useless since its already part of complementary solution. u need to make a guess thats not part of `complementary solution family`

OpenStudy (anonymous):

ohh.. i see, so thats why he kept adding x then? i noticed, is it right if i say, Axe^-x is identical to xe^-x, and (Ax^2+Bx)e^-x is identical (the B part which is Bxe^-x) that's why its useless?

ganeshie8 (ganeshie8):

yesss lets see why second guess is useless

ganeshie8 (ganeshie8):

first guess : \(Y_p = (Ax+B)e^{-x} \) second guess : \(Y_p = (Ax^2+ Bx)e^{-x} \)

ganeshie8 (ganeshie8):

how is the second guess part of complementary solution itself ?

OpenStudy (anonymous):

the Bxe^-x ? haha

ganeshie8 (ganeshie8):

yes, complementary solution : \(y_c(x) = c_1e^{-x} + c_2xe^{-x}\) plugin \(c_1 =A= 0, ~~ c_2 = B\) u get back ur second guess

ganeshie8 (ganeshie8):

so the second guess is bad too

OpenStudy (anonymous):

anyway going back to our previous question.. y"-y'-6y=1+e^-3x you put A because RHS is 1 (a constant) and you put Bxe^-3x so that it wont be the same as in the complementary solution?

ganeshie8 (ganeshie8):

for the original question, complementary solution : \(y_c(x) = c_1e^{-2x} + c_2e^{3x}\)

ganeshie8 (ganeshie8):

diff eq : \(y''-y'-6y=1+e^{-2x}\) our first guess : \(y_p = A + Be^{-2x}\)

ganeshie8 (ganeshie8):

clearly, when this is part of complementary solution family wid \(c_1 = B, ~c_2 = A = 0\)

ganeshie8 (ganeshie8):

its a bad guess

OpenStudy (anonymous):

i see now.. okay, how about for this: involving trig: 9y''+y=e^-3xsin(x/3)+ xcos(x/3) my guess would be, Ae^-3xsin(x/3)+ Be^-3xcos(x/3) what do you think?

OpenStudy (anonymous):

\[9y"+y=e ^{-3x}\sin (\frac{ x }{ 3 }) + xcos(\frac{ x }{ 3 })\]

ganeshie8 (ganeshie8):

complementary solution : \(y_c = c_2 \sin (\frac{x}{3}) + c_1\cos(\frac{x}{3}) \)

OpenStudy (anonymous):

oh it would be the same then, so means i need to add x somewhere

ganeshie8 (ganeshie8):

guess for sin(kt) : Asin(kt) + Bcos(kt)

ganeshie8 (ganeshie8):

I think u wil get 8 variables to solve

OpenStudy (anonymous):

woow why is that?

ganeshie8 (ganeshie8):

\(9y"+y=e ^{-3x}\sin (\frac{ x }{ 3 }) + x\cos(\frac{ x }{ 3 }) \) \(y_p=[Ae ^{-3x}] [B\sin (\frac{ x }{ 3 }) + C\cos(\frac{ x }{ 3 }) ] + [Dx+E][F\sin (\frac{ x }{ 3 }) + G\cos(\frac{ x }{ 3 })] \)

ganeshie8 (ganeshie8):

here is the particular solution from wolfram : http://prntscr.com/3e6kl4

ganeshie8 (ganeshie8):

doesnt look pleasant to solve so many constants ;/

OpenStudy (anonymous):

haha yeah xD no need to solve these just looking for Yp haha xD

OpenStudy (anonymous):

btw im not sure how u got your guess honestly

ganeshie8 (ganeshie8):

i used my table... ``` function : guess x : Ax + B e^(kx) : Ae^(kx) sin(kx) : Asin(kx) + Bcos(kx) cos(kx) : Asin(kx) + Bcos(kx) ```

OpenStudy (anonymous):

ooohh!! this useful! haha thank you! :D

ganeshie8 (ganeshie8):

np :)